# Qingdao, Shandong 266237, China

### v1

@@ Abstract
Deep Inelastic Scattering (Phenomenology), Spin-1 Target, Transverse Momentum Dependent Parton Distribution Functions
@@ 1 Introduction
Deep inelastic scattering (DIS) plays a significant role in studying the structure of the nucleon. The large momentum of the exchanged virtual photon, $q=l-l^{\prime}$ with $Q^{2}=-q^{2}$, provides the hard scale necessary to apply the factorization framework. In this case, the cross section for the inclusive DIS process can be approximately expressed as a convolution of the lepton-parton hard scattering and the collinear parton distribution functions (PDFs), $f_{q}(x)$, which indicates the probability density of finding a parton of flavor $q$ carrying a fraction $x$ of the nucleon momentum. Analogously to PDFs, fragmentation functions (FFs) are introduced to describe the probability density that a parton $q$ produces a hadron $h$ carrying a fraction $z$ of the parent parton. Inclusive DIS is characterized by a single hard scale and is not sensitive to the intrinsic motion of partons. To gain sensitivity to parton transverse momentum, we have to consider Semi-Inclusive DIS (SIDIS), i.e., DIS with an additional hadron detected in the final state. The transverse momentum $P_{h\perp}$ of the produced hadron provides an additional scale. In the region where $P_{h\perp}\ll Q$, Transverse-Momentum-Dependent (TMD) factorization can be applied Collins:1981uw ; Collins:1981uk ; Collins:2011zzd ; Rogers:2015sqa and the cross section can be described in terms of TMD PDFs, $f(x,\bm{p}_{T}^{2})$, and TMD FFs, $D(z,\bm{k}_{T}^{2})$, where $\bm{p}_{T}$ and $\bm{k}_{T}$ indicate the transverse momenta of the partons with respect to the nucleon and the produced hadron, respectively.
@@
Spin-1 hadrons and nuclei provide unique opportunities to explore novel effects that are inaccessible with spin-1/2 targets. Studies of the structure of spin-1 particles are still limited, although much progress has been made over the years Hoodbhoy:1988am ; Bora:1997pi ; Edelmann:1997qe ; Hino:1999qi ; Kumano:2020ijt ; Kumano:2024fpr ; Frankfurt:1983qs ; Hoodbhoy:1988am ; Chen:2016moq ; Kumano:2021fem ; Kumano:2020gfk ; Kumano:2019igu ; Kumano:2016ude ; Ninomiya:2017ggn ; Strikman:2017koc ; Kaur:2020emh ; Keller:2022abm ; Kumano:2021xau ; Sargsian:2022rmq ; Yang:2023zod ; Cosyn:2024drt ; Qiao:2024bgg ; Puhan:2023hio . Compared to spin-1/2 particles, spin-1 particles possess five additional independent tensor-polarized states, as described by their spin density matrix Bacchetta:2000jk . A complete set of TMD PDFs for a spin-1 particle at leading twist and higher has been derived from the decomposition of the quark-quark correlation function Bacchetta:2000jk ; Kumano:2020ijt ; Kumano:2024fpr , and can be used to describe both inclusive and semi-inclusive DIS processes with a spin-1 target. [^1] Similar studies for gluon-gluon correlation function were published in ref. Boer:2016xqr . The cross section for inclusive DIS off a spin-1 target can be expressed in terms of four tensor-polarized structure functions $b_{1-4}$ in addition to the ones for the spin-1/2 nucleons Frankfurt:1983qs ; Hoodbhoy:1988am . Among them, $b_{1}$ and $b_{2}$, as leading-twist structure functions, satisfying the Callan-Gross-like relation $xb_{1}=b_{2}$ in the Bjorken scaling limit Close:1990zw , where $x$ is the Bjorken variable. Although a sum rule $\int dx\,b_{1}(x)=0$ Close:1990zw can be derived when considering only the valence-quark contributions in the tensor structure, this does not imply that $b_{1}(x)=0$ for spin-1 hadrons and nuclei. A systematic analysis of jet production in SIDIS with a spin-1 target has been performed in ref. Chen:2020ugq .
@@
The deuteron is the simplest stable spin-1 target used in experiments and is a weakly bound system of a proton and a neutron. The so-called nuclear shadowing effect is considered to contribute significantly to the structure function $b_{1}$ at small $x$ Nikolaev:1996jy ; Bora:1997pi ; Edelmann:1997qe , and pions also play a crucial role in deuteron structure Miller:2013hla . The HERMES collaboration first measured $b_{1}$ via DIS off a tensor-polarized spin-1 deuteron and found that $b_{1}$ exhibits a steep rise as $x\rightarrow 0$ HERMES:2005pon . Subsequently, quark and antiquark distributions were extracted from a fit to the HERMES data Kumano:2010vz . Phenomenological studies based on these experimental data have been carried out in refs. Cosyn:2017fbo ; Cosyn:2020kwu ; Kumano:2019igu . The magnitude of $b_{1}$ and the $x$-dependence of $b_{1}$ calculated using the conventional convolution model exhibit a large difference from experimental results Cosyn:2017fbo ; Cosyn:2020kwu , which may suggest that a new hadron mechanism exists. A possible contribution from a hidden color state, a non-standard quark-gluon configuration, has been proposed to explain this discrepancy Miller:2013hla .
@@
Understanding tensor-polarized PDFs for spin-1 hadrons and nuclei can provide insights into their internal structure. Recent studies indicate that the calculation of deuteron PDFs on the lattice is very challenging Chen:2024rgi , which makes it necessary to extract the PDFs from experimental observables. The data points from the HERMES collaboration lie approximately within two standard deviations of zero, necessitating a measurement with higher precision. Recently, the CLAS12 collaboration has proposed to accurately measure $b_{1}$ by DIS off a tensor-polarized spin-1 deuteron at JLab Jlabproposal . The experiment could be extended to measure also tensor-polarized structure functions in the SIDIS process Poudel:2025tac ; Poudel:2025nof , thus making it possible to access tensor-polarized TMD PDFs. It has been also proposed to study the structure of spin-1 hadron in proton-deuteron Drell-Yan processes Hino:1998ww ; Hino:1999qi ; Kumano:1999bt ; Kumano:2016ude ; Qiao:2024bgg , which could be carried out by the SpinQuest experiment at Fermilab Keller:2022abm ; Keller:2020wan .
@@
In this paper, we present a symmetric study of the SIDIS process off a tensor-polarized spin-1 target. The differential cross section can be generally expressed in terms of 23 tensor-polarized structure functions, corresponding to specific angular distributions and target spin states. Within the kinematic region where the target and detected hadron are nearly back-to-back, we apply the TMD factorization to calculate the structure functions at first subleading twist accuracy (twist-3). We obtain 21 nonvanishing structure functions, which can be used to study the tensor-polarized TMD PDFs for spin-1 particles.
@@ 2 Kinematics and spin states
$x_{d}=\frac{Q^{2}}{2P\cdot q},\quad y=\frac{P\cdot q}{P\cdot l},\quad z=\frac{P\cdot P_{h}}{P\cdot q},\quad\gamma=\frac{2Mx_{d}}{Q},$ (2)
where $x_{d}$ is the scaling variable for the deuteron. The Bjorken scaling variable, defined as $x_{B}=Q^{2}/(2M_{N}\nu)$ with the nucleon mass $M_{N}$, is related to $x_{d}$ via $x_{B}=2x_{d}$. Following the Trento conventions Bacchetta:2004jz , we choose the virtual photon-target frame, as illustrated in figure 1. The azimuthal angle $\phi_{h}$ spanned by the lepton plane and the hadron plane can be expressed in a Lorentz invariant form as
@@
with nonzero components $g_{\perp}^{11}=g_{\perp}^{22}=-1$ and $\epsilon_{\perp}^{12}=-\epsilon_{\perp}^{21}=-1$, where the convention for totally antisymmetric tensor is $\epsilon^{0123}=1$.
As defined in ref. Bacchetta:2000jk , the polarization for spin-1 particles is characterized by both the spin vector $S^{\mu}$ and the spin tensor $T^{\mu\nu}$. We can decompose these spin tensors into longitudinal and transverse components with respect to the momentum of the hadron. In light-cone coordinates, the hadron momentum can be written as
@@
The term $a^{\{\mu}b^{\nu\}}$ indicates the symmetrization of the indices, i.e., $a^{\{\mu}b^{\nu\}}=a^{\mu}b^{\nu}+a^{\nu}b^{\mu}$. The four-vector $S_{LT}^{\mu}$ and tensor $S_{TT}^{\mu\nu}$ refer to the longitudinal-transverse and transverse-transverse tensor polarization. The corresponding systematic definitions and physical interpretations have been discussed in ref. Bacchetta:2000jk .
@@
3 The cross section in terms of structure functions
@@
contributing to the antisymmetric part of $W^{\mu\nu}$, where the subscript $T$ indicates the tensor-polarized states.
@@
where $d\sigma_{U}$, $d\sigma_{V}$, and $d\sigma_{T}$ represent the contributions from the unpolarized, vector-polarized, and tensor-polarized states, respectively. To emphasize the tensor-polarized dependent cross section, we denote it with the subscript $\rm Tens$” in the following. According to the azimuthal modulations and the spin states of the target, the cross section can be expressed in terms of 41 structure functions. Among them, the 23 tensor-polarized ones only exist when the target has spin $s\geq 1$, which is also included in ref. Chen:2020ugq , while the remaining 18 ones also exist for SIDIS with a spin-1/2 target, as defined in ref. Bacchetta:2006tn . Here we give the explicit expression of the tensor-polarized dependent cross section as
@@
$\sum_{h}\int dzz\frac{d\sigma(\ell N\rightarrow\ell hX)}{dzdx_{d}dyd\psi}=\frac{1}{\nu}\sum_{h}\int dE_{h}\,E_{h}\frac{d\sigma(\ell N\rightarrow\ell hX)}{dE_{h}dx_{d}dyd\psi}=\frac{\nu+M}{\nu}\frac{d\sigma(\ell N\rightarrow\ell X)}{dx_{d}dyd\psi},$ (44)
where $E_{h}=(P\cdot P_{h})/M$ and $\nu=(P\cdot q)/M$ are the energies of the produced hadron and the virtual photon, respectively. When summing over all hadrons in the final state, the total energy is given by $\nu+M$. With the relation $(\nu+M)/\nu=1+\gamma^{2}/(2x_{d})$, we have
@@
$F_{U(LL),T}$ $=-\biggl{[}2\bigl{(}1+\gamma^{2}\bigr{)}b_{1}-\frac{\gamma^{2}}{x_{d}}\biggl{(}\frac{1}{6}b_{2}-\frac{1}{2}b_{3}\biggr{)}\biggr{]},$ (48)
$F_{U(LL),L}$ $=\frac{1}{x_{d}}\biggl{[}2\bigl{(}1+\gamma^{2}\bigr{)}x_{d}b_{1}-(1+\gamma^{2})^{2}\biggl{(}\frac{1}{3}b_{2}+b_{3}+b_{4}\biggr{)}$
@@
$F_{U(LT)}^{\cos\phi_{LT}}$ $=-\frac{\gamma}{2x_{d}}\biggl{[}\bigl{(}1+\gamma^{2}\bigr{)}\biggl{(}\frac{1}{3}b_{2}-b_{4}\biggr{)}+\biggl{(}\frac{2}{3}b_{2}-2b_{3}\biggr{)}\biggr{]},$ (50)
@@
The structure functions $b_{1}$ and $b_{2}$ obey a Callan-Gross-like relation $b_{2}=2x_{d}b_{1}$ in the Bjorken limit. The structure functions defined in eq. (45) have advantages in illustrating the geometric dependence on the different tensor-polarized states and azimuthal modulations.
4 The structure functions in terms of partonic functions
4.1 The hadronic tensor up to subleading order in $1/Q$
@@
Here, we label the two light-cone vectors by $n_{+}$ and $n_{-}$ to distinguish them with the definitions of $\bar{n}$ and $n$ in eq. (12). They can be expressed in terms of $\hat{t}$, $\hat{z}$, and the transverse momentum $q_{T}$ Mulders:1995dh ; Boer:2003cm ,
@@
with the light-cone longitudinal momentum fraction defined by $xP^{+}=P_{h}^{-}/z=Q/\sqrt{2}$. $p_{T}$ and $k_{T}$ correspond to the transverse momenta of the quark distributing in the target and the quark fragmenting into the observed hadron, respectively. The sum over “aa” runs over all active quark and antiquark flavors, and $e_{a}$ denotes the charge in units of the positron charge. The terms with $n_{+}$ and $n_{-}$ in eq. (55) arise from fermion propagators in the quark-photon hard scattering part while neglecting the suppression by $Q^{2}$. The explicit forms of the quark distribution correlator $\Phi$, the quark fragmentation correlator $\Delta$, and the quark-gluon-quark correlators $\tilde{\Phi}_{A}$ and $\tilde{\Delta}_{A}$ with an additional gluon leg will be provided in the following subsections. In eq. (55), the first, second, and third terms correspond to the diagrams shown in figures 2(a)2(b), and 2(c), where the gluon is transverse polarized. The h.c. in eq. (25) stands for the Hermitian conjugate of figures 2(b) and 2(c), with the gluon on the opposite side of the final-state cut.
4.2 The quark-quark correlators

### v2

@@ Abstract
Deep Inelastic Scattering or Small-x Physics, Parton Distributions
arxiv: 2508.06134
@@ 1 Introduction
Deep inelastic scattering (DIS) plays a significant role in studying the structure of the nucleon. The large momentum of the exchanged virtual photon, $q$ with $Q^{2}=-q^{2}$, provides the hard scale necessary to apply the factorization framework. In this case, the cross section for the inclusive DIS process can be approximately expressed as a convolution of the lepton-parton hard scattering and the collinear parton distribution functions (PDFs), $f_{q}(x)$, which indicates the probability density of finding a parton of flavor $q$ carrying a fraction $x$ of the nucleon momentum. Analogously to PDFs, fragmentation functions (FFs) are introduced to describe the probability density that a parton $q$ produces a hadron $h$ carrying a fraction $z$ of the parent parton. Inclusive DIS is characterized by a single hard scale and is not sensitive to the intrinsic motion of partons. To gain sensitivity to parton transverse momentum, we have to consider Semi-Inclusive DIS (SIDIS), i.e., DIS with an additional hadron detected in the final state. The transverse momentum $P_{h\perp}$ of the produced hadron provides an additional scale. In the region where $P_{h\perp}\ll Q$, Transverse-Momentum-Dependent (TMD) factorization can be applied Collins:1981uw ; Collins:1981uk ; Collins:2011zzd ; Rogers:2015sqa and the cross section can be described in terms of TMD PDFs, $f(x,\bm{p}_{T}^{2})$, and TMD FFs, $D(z,\bm{k}_{T}^{2})$, where $\bm{p}_{T}$ and $\bm{k}_{T}$ indicate the transverse momenta of the partons with respect to the nucleon and the produced hadron, respectively.
@@
Spin-1 hadrons and nuclei provide unique opportunities to explore novel effects that are inaccessible with spin-1/2 targets. Studies of the structure of spin-1 particles are still limited, although much progress has been made over the years Bora:1997pi ; Edelmann:1997qe ; Hino:1999qi ; Kumano:2020ijt ; Kumano:2024fpr ; Frankfurt:1983qs ; Hoodbhoy:1988am ; Chen:2016moq ; Kumano:2021fem ; Kumano:2020gfk ; Kumano:2019igu ; Kumano:2016ude ; Ninomiya:2017ggn ; Strikman:2017koc ; Kaur:2020emh ; Keller:2022abm ; Kumano:2021xau ; Sargsian:2022rmq ; Yang:2023zod ; Cosyn:2024drt ; Qiao:2024bgg ; Puhan:2023hio . Compared to spin-1/2 particles, spin-1 particles possess five additional independent tensor-polarized components, as described by their spin density matrix Bacchetta:2000jk . A complete set of TMD PDFs for a spin-1 particle at leading twist and higher has been derived from the decomposition of the quark-quark correlation function Bacchetta:2000jk ; Kumano:2020ijt ; Kumano:2024fpr , and can be used to describe both inclusive and semi-inclusive DIS processes with a spin-1 target. [^1] Similar studies for gluon-gluon correlation function were published in ref. Boer:2016xqr . The cross section for inclusive DIS off a spin-1 target can be expressed in terms of four tensor-polarized structure functions $b_{1-4}$ in addition to the ones for the spin-1/2 nucleons Frankfurt:1983qs ; Hoodbhoy:1988am . Among them, $b_{1}$ and $b_{2}$, as leading-twist structure functions, satisfying the Callan-Gross-like relation $xb_{1}=b_{2}$ in the Bjorken scaling limit Close:1990zw , where $x$ is the Bjorken variable. Although a sum rule $\int dx\,b_{1}(x)=0$ Close:1990zw can be derived when considering only the valence-quark contributions in the tensor structure, this does not imply that $b_{1}(x)=0$ for spin-1 hadrons and nuclei. A systematic analysis of jet production in SIDIS with a spin-1 target has been performed in ref. Chen:2020ugq .
@@
The deuteron is the simplest stable spin-1 target used in experiments and is a weakly bound system of a proton and a neutron. The so-called nuclear shadowing effect is considered to contribute significantly to the structure function $b_{1}$ at small $x$ Nikolaev:1996jy ; Bora:1997pi ; Edelmann:1997qe , and pions also play a crucial role in deuteron structure Miller:2013hla . The HERMES collaboration first measured $b_{1}$ via DIS off a tensor-polarized spin-1 deuteron and found that $b_{1}$ exhibits a steep rise as $x\rightarrow 0$ HERMES:2005pon . Subsequently, quark and antiquark distributions were extracted from a fit to the HERMES data Kumano:2010vz . Phenomenological studies based on these experimental data have been carried out in refs. Cosyn:2017fbo ; Cosyn:2020kwu ; Kumano:2019igu . The magnitude of $b_{1}$ and the $x$-dependence of $b_{1}$ calculated using the conventional convolution model exhibit a large difference from experimental results Cosyn:2017fbo , which may suggest that a new hadron mechanism exists. A possible contribution from a hidden color state, a non-standard quark-gluon configuration, has been proposed to explain this discrepancy Miller:2013hla .
@@
Understanding tensor-polarized PDFs for spin-1 hadrons and nuclei can provide insights into their internal structure. Recent studies indicate that the calculation of deuteron PDFs on the lattice is very challenging Chen:2024rgi , which makes it necessary to extract the PDFs from experimental observables. The data points from the HERMES collaboration lie approximately within two standard deviations of zero, necessitating a measurement with higher precision. Experiments have been proposed at JLab to measure tensor-polarized structure functions in the inclusive and semi-inclusive DIS processes Jlabproposal ; Poudel:2025tac ; Poudel:2025nof , thus making it possible to access tensor-polarized TMD PDFs. It has been also proposed to study the structure of spin-1 hadron in proton-deuteron Drell-Yan processes Hino:1998ww ; Hino:1999qi ; Kumano:1999bt ; Kumano:2016ude ; Qiao:2024bgg , which could be carried out by the SpinQuest experiment at Fermilab Keller:2022abm ; Keller:2020wan .
@@
In this paper, we present a systematic study of the SIDIS process off a tensor-polarized spin-1 target. The differential cross section can be generally expressed in terms of 23 tensor-polarized structure functions, corresponding to specific angular distributions and target spin states. Within the kinematic region where the target and detected hadron are nearly back-to-back, we apply the TMD factorization to calculate the structure functions at first subleading twist accuracy (twist-3). We obtain 21 nonvanishing structure functions, which can be used to study the tensor-polarized TMD PDFs for spin-1 particles.
@@ 2 Kinematics and spin states
$x_{d}=\frac{Q^{2}}{2P\cdot q},\quad y=\frac{P\cdot q}{P\cdot l},\quad z=\frac{P\cdot P_{h}}{P\cdot q},\quad\gamma=\frac{2Mx_{d}}{Q},\quad\nu=\frac{P\cdot q}{M},$ (2)
where $x_{d}$ is the scaling variable for the deuteron. The Bjorken scaling variable, defined as $x_{B}=Q^{2}/(2M_{N}\nu)$ with the nucleon mass $M_{N}$, is related to $x_{d}$ via $x_{B}\approx 2x_{d}$. Following the Trento conventions Bacchetta:2004jz , we choose the virtual photon-target frame, as illustrated in figure 1. The azimuthal angle $\phi_{h}$ spanned by the lepton plane and the hadron plane can be expressed in a Lorentz invariant form as
@@
which have nonzero components $g_{\perp}^{11}=g_{\perp}^{22}=-1$ and $\epsilon_{\perp}^{12}=-\epsilon_{\perp}^{21}=-1$ in the virtual photon-target frame, where the convention for totally antisymmetric tensor is $\epsilon^{0123}=1$.
As defined in ref. Bacchetta:2000jk , the polarization for spin-1 particles is characterized by both the spin vector $S^{\mu}$ and the spin tensor $T^{\mu\nu}$. We define the transverse and longitudinal spin components with respect to the three-momentum of the hadron. In light-cone coordinates, the hadron momentum can be written as
@@
When $n$ and $\bar{n}$ are chosen as in eq. (12), one can obtain the expression of $g_{\perp}^{\mu\nu}$ in eq. (4). The term $a^{\{\mu}b^{\nu\}}$ indicates the symmetrization of the indices, i.e., $a^{\{\mu}b^{\nu\}}=a^{\mu}b^{\nu}+a^{\nu}b^{\mu}$. The four-vector $S_{LT}^{\mu}$ and tensor $S_{TT}^{\mu\nu}$ refer to the longitudinal-transverse and transverse-transverse tensor polarization. The corresponding systematic definitions and physical interpretations have been discussed in ref. Bacchetta:2000jk .
@@
3 Cross section in terms of structure functions
@@
contributing to the antisymmetric part of $W^{\mu\nu}$, where the subscript $T$ indicates the tensor-polarized components.
@@
where $d\sigma_{U}$, $d\sigma_{V}$, and $d\sigma_{T}$ represent the contributions from the unpolarized, vector-polarized, and tensor-polarized components, respectively. To emphasize the tensor-polarized dependent cross section, we denote it with the subscript $\rm Tens$” in the following. According to the azimuthal modulations and the spin states of the target, the cross section can be expressed in terms of 41 structure functions. Among them, the 23 tensor-polarized ones only exist when the target has spin $s\geq 1$, which is also included in ref. Chen:2020ugq , while the remaining 18 ones also exist for SIDIS with a spin-1/2 target, as defined in ref. Bacchetta:2006tn . Here we give the explicit expression of the tensor-polarized dependent cross section as
@@
$\sum_{h}\int dzz\frac{d\sigma(\ell d\rightarrow\ell hX)}{dzdx_{d}dyd\psi}=\frac{1}{\nu}\sum_{h}\int dE_{h}\,E_{h}\frac{d\sigma(\ell d\rightarrow\ell hX)}{dE_{h}dx_{d}dyd\psi}=\frac{\nu+M}{\nu}\frac{d\sigma(\ell d\rightarrow\ell X)}{dx_{d}dyd\psi},$ (44)
where $E_{h}=(P\cdot P_{h})/M$ is the energy of the produced hadron in the target rest frame. When summing over all hadrons in the final state, the total energy is given by $\nu+M$. Factoring out the overall kinematic factor, $(\nu+M)/\nu=1+\gamma^{2}/(2x_{d})$, originated from the integration over $z$ and the sum over hadrons, we have
@@
$F_{U(LL),T}$ $=-\frac{2}{3}\biggl[2\bigl(1+\gamma^{2}\bigr)b_{1}-\frac{\gamma^{2}}{x_{d}}\biggl(\frac{1}{6}b_{2}-\frac{1}{2}b_{3}\biggr)\biggr],$ (48)
$F_{U(LL),L}$ $=\frac{2}{3x_{d}}\biggl[2\bigl(1+\gamma^{2}\bigr)x_{d}b_{1}-(1+\gamma^{2})^{2}\biggl(\frac{1}{3}b_{2}+b_{3}+b_{4}\biggr)$
@@
$F_{U(LT)}^{\cos\phi_{LT}}$ $=-\frac{\gamma}{4x_{d}}\biggl[\bigl(1+\gamma^{2}\bigr)\biggl(\frac{1}{3}b_{2}-b_{4}\biggr)+\biggl(\frac{2}{3}b_{2}-2b_{3}\biggr)\biggr],$ (50)
@@
The structure functions $b_{1}$ and $b_{2}$ obey a Callan-Gross-like relation $b_{2}=2x_{d}b_{1}$ in the Bjorken limit. The structure functions defined in eq. (45) have advantages in illustrating the geometric dependence on the different tensor-polarized components and azimuthal modulations. We note that our conventions of the tensor polarization components differ from those used in ref. Cosyn:2024drt . They are related by $T_{LL}=2S_{LL}/3$, $T_{LT}\cos\phi_{T_{L}}=S_{LT}^{x}/2$, and $T_{TT}\cos 2\phi_{T_{T}}=S_{TT}^{xx}$, where $T_{LL}$, $_{LT}, TT$$_{TT}$, $\phi_{T_{L}}$, and $\phi_{T_{T}}$ are the tensor polarization components and azimuthal angles used in ref. Cosyn:2024drt . These differences can be absorbed into the normalization of corresponding structure functions, and the formal expression of differential cross section keeps the same. Accordingly, the right-hand sides of eqs. (48) and (49) differ by a factor of $2/3$ from eqs. (17a) and (17b) in ref. Cosyn:2024drt , and the right-hand side of eq. (50) differs by a factor of $1/2$ from eq. (17d) in ref. Cosyn:2024drt . The corresponding factors due to such normalization differences are also reflected in eqs. (164)-(167).
4 Structure functions in terms of partonic functions
4.1 Hadronic tensor up to subleading order in $1/Q$
@@
Here, we label the two light-cone vectors by $n_{+}$ and $n_{-}$ to distinguish them from the definitions of $\bar{n}$ and $n$ in eq. (12). They can be expressed in terms of $\hat{t}$, $\hat{z}$, and the transverse momentum $q_{T}$ Mulders:1995dh ; Boer:2003cm ,
@@
with the light-cone longitudinal momentum fraction defined by $xP^{+}=P_{h}^{-}/z=Q/\sqrt{2}$. Here $x$ equals the scaling variable $x_{d}$ at the leading order when neglecting the target mass. The $p_{T}$ and $k_{T}$ correspond to the transverse momenta of the quark distributing in the target and the quark fragmenting into the observed hadron, respectively. The sum over “aa” runs over all active quark and antiquark flavors, and $e_{a}$ denotes the charge in units of the positron charge. The terms with $n_{+}$ and $n_{-}$ in eq. (55) arise from fermion propagators in the quark-photon hard scattering part while neglecting the suppression by $Q^{2}$. The explicit forms of the quark distribution correlator $\Phi$, the quark fragmentation correlator $\Delta$, and the quark-gluon-quark correlators $\tilde{\Phi}_{A}$ and $\tilde{\Delta}_{A}$ with an additional gluon leg will be provided in the following subsections. In eq. (55), the first, second, and third terms correspond to the diagrams shown in figures 2(a)2(b), and 2(c), where the gluon is transverse polarized. The h.c. in eq. (25) stands for the Hermitian conjugate of figures 2(b) and 2(c), with the gluon on the opposite side of the final-state cut.
4.2 Quark-quark correlators